# Reinforcement Learning and Machine ethics: a systematic review

### v1

@@ Abstract
Machine ethics is the field that studies how ethical behaviour can be accomplished by autonomous systems. While there exist some systematic reviews aiming to consolidate the state of the art in machine ethics prior to 2020, these tend to not include work that uses reinforcement learning agents as entities whose ethical behaviour is to be achieved. The reason for this is that only in the last years we have witnessed an increase in machine ethics studies within reinforcement learning. We present here a systematic review of reinforcement learning for machine ethics and machine ethics within reinforcement learning. Additionally, we highlight trends in terms of ethics specifications, components and frameworks of reinforcement learning, and environments used to result in ethical behaviour. Our systematic review aims to consolidate the work in machine ethics and reinforcement learning thus completing the gap in the state of the art machine ethics landscape.
[…]
1414
@@ 1 Introduction
Machine ethics “is concerned with the behaviour of machines towards human users and other machines” [8]. This includes studying how to implement machine reasoning in a way that includes the consideration of ethical factors [9, 75]. Since a 2006 AAAI Workshop on Machine Ethics [8], there has been a rapid expansion in proposed systems and formalisms for performing machine ethics (e.g,  [73, 28, 19, 17, 40, 30]) as evident from surveys of the field [44, 71, 77].
@@
Reviews of machine ethics, at present, cover the period up to 2020. Currently, we observe a notable trend in reinforcement learning, specifically focusing on the ethical behavior of reinforcement learning agents.
@@
In reinforcement learning (RL) [68] an artificial agent makes decisions in an environment and receives rewards and punishments as outcomes of those decisions. The goal, typically of the learning process, is to find a policy on how to make the decisions so that an optimal reward is attained. Reinforcement learning, and specifically deep reinforcement learning that combines neural networks and reinforcement learning, has been of particular recent interest since its successful application in training artificial agents to play human level computer games [42]. As reinforcement learning is about training an artificial agent how to make decisions, it is very intuitive as to why ethical decision-making would receive attention in the reinforcement learning community.
@@
Early papers on the reinforcement learning and machine ethics topic, such as [11], gained recognition, along with subsequent work by researchers  [13, 34, 43, 56, 45, 70, 74]. There is a gap in the understanding of the state of the art in machine learning if that understanding does not include the most recent work on reinforcement learning. Our goal is to bridge this gap by conducting a literature review on reinforcement learning in the context of machine ethics.
@@
Following a keywords search in the usual repositories of artificial intelligence research articles we found 605 potential papers which we have narrowed down to 62 papers that were on topic and which were carefully analysed. We analysed the papers based on the RL paradigm they deploy, how they modify which RL “component" to achieve ethical behaviour, and which ethical theory they directly or indirectly implement as well as what kind of implementation examples the papers use. Beyond this we observe and discuss the trends in this body of work and draw recommendation for future research in this field.
@@
This article contributes to both the disciplines of reinforcement learning and machine ethics by establishing a clear state of the art of the intersection of these two fields and giving future work a good anchor to position and grow from existing experience. Furthermore, we highlight challenges and possible pitfalls of reinforcement learning and machine ethics, thus ensuring a responsible and sustainable development of what at present appears to be a trendy and growing research field.
Background and related work
@@
In this section, we discuss reinforcement learning and the various algorithms employed in the papers we review. This is followed by a discussion on previous works in machine ethics and relevant gaps.
Reinforcement Learning
@@
Reinforcement learning (RL) is a machine learning paradigm in which an agent navigates an environment with states $s\in S$, actions $a\in A$, next states $s^{\prime}\in S^{\prime}$ and rewards $r\in R$. Such an agent learns through trial and error an optimum behavior which maximizes a numerical reward function [68]. This optimum behavior is denoted by $\pi^{*}$, or the optimum policy. Markov decision processes (MDP) are a mathematically idealized form of the RL problem. In an MDP, the probability of each possible value of the next state $S^{\prime}$ and reward $R$ depends on the immediately preceding state and action, $S$ and $A$, completely characterize the environment and that these states, actions and rewards are finite.
@@
A method to solve MDPs is by using value function approximation, where a numerical value is calculated for being in a state $V(s)$ and committing an action in a given state $Q(s,a)$. As the agent visits different states in the environment, a value table is updated based on the reward obtained. This method is often referred to as Q-learning, where a Q table is updated and based on these Q-values an optimal policy is determined. However, an issue with such an approach is that in an environment with a very high state space, such as the game of chess, it can be very expensive to compute a Q-table.
@@
In high state-space scenario, a neural network is often used to encode the state space and output the Q-value for being in that state. This paradigm of using a neural network in the Q-learning algorithm is referred to as Deep Q-learning. There are, however, limitations with using value function approximation, since there is an exploration-exploitation conundrum. These networks are said to be more exploratory in nature and often fail to converge to an optimum. To overcome this issue, actor-critic methods are used where two neural networks, the actor and the critic are used to predict action probabilities and Q-values, respectively, instead of only calculating Q-values. Methods such as PPO, DDPG etc are popular actor-critic methods.
@@
There are other types of RL, such inverse RL, multi-objective RL and co-operative RL. In inverse RL (IRL) [1], the agent learns the reward functions in an environment, given the policy. This is useful in applications such as robotics, where the robot can "watch" a human demonstrate a task (optimal policy) and then the robot can imitate the same, while learning a reward function in its environment. Next, multi-objective RL (MORL), an agent is trained to optimize conflicting objective functions, demonstrated in works by Rodriguez et al. [59] and servral others discussed in this review. Finally, co-operative inverse reinforcement learning, is a variant of IRL in a multi-agent scenario. However, unlike classical IRL, where the human is the basis for learning, the other agents in the system often perform active teaching, learning and communicative actions [33].
Systematic reviews in machine ethics
@@
We here briefly summarise the findings of the three reviews in machine ethics, their findings and methodology.
@@
Tolmeijer et al. [71] focused on the implementations of machine ethics specifically and found 49 papers from Web of Science, Scopus, ACM Digital Library, Wiley Online Library, ScienceDirect, AAAI Publications, Springer Link, and IEEE Xplore. These papers were then analysed based on types of ethical theory they implemented, the nontechnical aspects when implementing those theories, and technological details.
@@
Nallur [44] considers the landscape of machine ethics implementations focusing on the different problems, techniques, methodologies and evaluation methodologies that are used. He did not do a systematic review, but rather a more qualitative analysis of selected prominent works.
@@
Yu et al [77] survey advances in techniques for incorporating ethics into AI. They consider publications in “leading AI research conferences including AAAI, AAMAS, ECAI and IJCAI, as well as articles from well-known journals”. They organise the papers into four categories: Exploring Ethical Dilemmas, Individual Ethical Decision Frameworks, Collective Ethical Decision Frameworks; and 4. Ethics in Human-AI Interactions. There is of course significant overlap among the papers included in all three surveys.
@@
We here follow to a large degree the approach of Tolmeijer et al. [71] in terms of where we search for papers. This is because Tolmeijer et al. [71] is the only systematic review in machine ethics. Understanding the state of the art in machine ethics is not without challenges. The papers published in this area are sent to a wide spectrum of venues. Machine ethics is also not a very large area of research and it is not very precisely positioned in the context of the AI alignment and AI Ethics initiatives. Therefore, it has to be accepted that any systematic review of the field is likely to not be exhaustive, even if we limit ourselves to a specific domain or methodology such as reinforcement learning.
@@ 2 Survey methodology
In this section, we describe the strategy used to find papers related to machine ethics and reinforcement learning, following which we discuss the process used to review the found articles.
Search strategy
Our review on implementations in machine ethics focuses on RL, hence we mandated this terminology while querying the databases. Since RL is to be used towards machine ethics implementations, we used alternative search terms such as artificial morality [5], machine morality and robot ethics [41], computational ethics [10], roboethics [26], artificial moral agents [22], and value alignment [59]. These terms have been used interchangeably to communicate the embedding of ethical theories into machines/computers/artificial intelligence agents. We show the query used below:
"reinforcement learning" AND ( "machine ethics" OR "artificial morality" OR "machine morality" OR "computational ethics" OR "roboethics" OR "robot ethics" OR "artificial moral agents" OR "value alignment")
The query was used to search the following databases 1. In the right hand side column we give the number of ‘hits’ from the query. Further filtering was performed on these ‘’hits.
@@
ACM Digital Library 158
IEEE Xplore 51
ScienceDirect 21
Scopus 41
SpringerLink 146
Web of Science 16
Willey Online Library 172
Total 605
@@
Selection process
A total of 605 papers were further filtered based on relevance to RL and machine ethics (ME). We included only journal articles, conference papers, and book chapters which were published before 4th January 2024. The diagram on Figure 1 shows the elimination process we used to arrive at the final list of manuscripts.
Figure 1: Selection process used to eliminate irrelevant articles but keep the relevant ones. First is to check whether the article focuses on RL, followed by checking if the RL algorithm is used for moral decision making, and finally to eliminate duplicates that appeared across multiple databases.
After applying this selection criteria, we ended up with 62 papers for review. Subsequently, we conducted a more focused review by filtering the papers based on their relevance to both machine ethics and RL, resulting in a total of 48 papers. In the next section, we delve deeper into the taxonomy creation and review process.
The analysis
In this section, we discuss the taxonomy and categorizations used to organize our survey. Since the focus is on RL, we closely examined its components such as the reward, policy, and value function. Additionally, we considered the specific RL paradigms used, allowing us to observe trends within the research community. Furthermore, we explore ethical theories and implementation examples, which constitute the core of machine ethics implementations. Finally, we categorize trends in machine ethics for RL based on contribution type, whether they are argumentative, empirical, mathematical, or a combination.
Component of RL Modified
There are several aspects that affect RL algorithms and their performance on tasks. For instance, the challenge of designing a good reward function to correctly incentivize agents can make or break the agent’s performance. Other aspects such as policy functions, value functions, etc., also play a part. In terms of developing RL algorithms which make more ethical choices, several researchers modify either one or more of these components to achieve the desired result. We categorized the research articles based on which component of RL was modified and we present our findings in Table 2.

### v2

@@ Abstract
Machine ethics is the field that studies how ethical behaviour can be accomplished by autonomous systems. While there exist some systematic reviews aiming to consolidate the state of the art in machine ethics prior to 2020, these tend to not include work that uses reinforcement learning agents as entities whose ethical behaviour is to be achieved. The reason for this is that only in the last years have we witnessed an increase in machine ethics studies within reinforcement learning. We present here a systematic review of reinforcement learning for machine ethics and machine ethics within reinforcement learning. Additionally, we highlight trends in terms of ethics specifications, components and frameworks of reinforcement learning, and environments used to result in ethical behaviour. Our systematic review aims to consolidate the work in machine ethics and reinforcement learning, thus filling the gap in the state of the art machine ethics landscape.
keywords:
Machine Ethics, Reinforcement Learning, Moral Philosophy
@@ 1 Introduction
Machine ethics “is concerned with the behaviour of machines towards human users and other machines” [1]. This includes studying how to implement machine reasoning in a way that includes the consideration of ethical factors [2, 3]. Since the 2006 AAAI Workshop on Machine Ethics [1], there has been a rapid expansion in proposed systems and formalisms for performing machine ethics (e.g,  [4, 5, 6, 7, 8, 9]) as evident from surveys of the field [10, 11, 12].
@@
Reviews of machine ethics, at present, cover the period up to 2020. Currently, we observe a notable trend in reinforcement learning, specifically focusing on the ethical behaviour of reinforcement learning agents.
@@
In reinforcement learning (RL) [13] an artificial agent makes decisions in an environment and receives rewards and punishments as outcomes of those decisions. The goal, typically of the learning process, is to find a policy on how to make decisions so that an optimal reward is attained. RL, and specifically deep RL that combines neural networks and RL, has been of particular recent interest since its successful application in training artificial agents to play human-level computer games [14]. Since RL is about training an artificial agent in how to make decisions; it is very intuitive as to why ethical decision-making would receive attention in the RL community.
@@
Early papers on the RL and machine ethics topic, such as [15], gained recognition, along with subsequent work by researchers  [16, 17, 18, 19, 20, 21, 22]. There is a gap in the understanding of the state of the art in machine ethics if that understanding does not include the most recent work on RL. Our goal is to bridge this gap by conducting a literature review on RL in the context of machine ethics. While our motivation derives from the observation that much work in this sub-field has taken place since the survey articles of 2020, we included earlier works in our survey in order to present a complete view of the application of RL to machine ethics, including important early results.
We used a systematic review approach similar to that outlined in [23]. We defined a review protocol that specifies a research question and the methods used to perform the review; we defined and documented our search strategy; we defined the inclusion and exclusion criteria; and we specified the information to be obtained from each paper identified by the process.
@@
Following a keywords search in the usual repositories of artificial intelligence research articles, we found 808 potential papers which we have narrowed down to 71 papers that were on topic and which were carefully analysed. We analysed the papers based on the RL framework they deploy, how they modify which RL “component” to achieve ethical behaviour, and which ethical theory they directly or indirectly implement, as well as what kind of implementation examples the papers use. Beyond this, we observe and discuss the trends in this body of work and draw recommendations for future research in this field.
@@
This article contributes to both the disciplines of RL and machine ethics by establishing a clear state of the art of the intersection of these two fields and giving future work a good anchor to position and grow from existing experience. Furthermore, we highlight the challenges and possible pitfalls of RL and machine ethics, thus ensuring a responsible and sustainable development of what at present appears to be a trendy and growing research field.
2 Background and related work
@@
In this section, we discuss RL and the various algorithms employed in the papers we review. This is followed by a discussion on previous work in machine ethics and relevant gaps.
2.1 Reinforcement Learning
@@
Reinforcement learning (RL) is a machine learning framework in which an agent navigates an environment with states $s\in S$, actions $a\in A$, next states $s^{\prime}\in S^{\prime}$ and rewards $r\in R$. Such an agent learns through trial and error an optimum behaviour which maximises a numerical reward function [13]. This optimum behaviour is denoted by $\pi^{*}$, or the optimum policy. Markov decision processes (MDP) are a mathematically idealised form of the RL problem. In an MDP, the probability of each possible value of the next state $s^{\prime}$ and reward $r$ depends on the immediately preceding state and action, $s$ and $a$, completely characterise the environment and that these states, actions, and rewards are finite. An important conundrum in RL is the exploration-exploitation trade-off, where if the agent performs over-exploration, the agent may not converge to an optimum, while if the agent over-exploits, the agent could potentially waste time on poor actions and fail to consider better alternatives [13].
@@
A method to solve MDPs is using value function approximation, where a numerical value is calculated for being in a state $V(s)$ and committing an action in a given state $Q(s,a)$. As the agent visits different states in the environment, a value table is updated based on the reward obtained. This method is often referred to as Q-learning, where a Q table is updated and based on these Q-values an optimal policy is determined. However, an issue with such an approach is that in an environment with a very high state space, such as the game of chess, it can be very expensive to compute a Q-table.
@@
In high state-space scenarios, a neural network is often used to encode the state space and output the Q-value for being in that state. This framework of using a neural network in the Q-learning algorithm is referred to as Deep Q-learning. There are, however, limitations to using value function approximation, since there is an exploration-exploitation conundrum. These networks are said to be more exploratory in nature and often fail to converge to an optimum [13]. Such behaviour can be exacerbated in environments with high stochasticity or vast state spaces, where the balance between continual exploration and fine-tuning of previously learned policies becomes increasingly challenging. To overcome this issue, actor-critic methods are used where two neural networks, an actor and a critic, are used to predict action probabilities and Q-values, respectively, instead of only calculating Q-values. Methods such as Proximal Policy Optimization (PPO), Deep Deterministic Policy Gradient (DDPG), etc. are popular actor-critic methods.
@@
There are other types of RL, such as inverse RL, multi-objective RL and co-operative RL. In inverse RL (IRL) [24], the agent learns the reward functions in an environment, given the policy. This is useful in applications such as robotics, where the robot can ”watch” a human demonstrate a task (optimal policy) and then the robot can imitate the same, while learning a reward function in its environment. Next, multi-objective RL (MORL), an agent is trained to optimise conflicting objective functions, demonstrated in works by Rodriguez et al. [25] and several others discussed in this review. Finally, co-operative inverse RL is a variant of IRL in a multi-agent scenario. However, unlike classical IRL, where the human is the basis for learning, the other agents in the system often perform active teaching, learning and communicative actions [26]. Other RL frameworks such as safe/constrained RL are often employed to prevent hazardous failures in RL algorithms during training and deployment [27]. While both terminologies are sometimes used interchangeably, constrained RL is a subset of safe RL that formally defines constraints and maximises rewards while meeting constraints.
2.2 Moral philosophy
Ethics being the crucial piece of the machine ethics puzzle, is succinctly summarised here. The three major ethical theories in moral philosophy are: deontology, consequentialist and virtue ethics. A consequentialist [11] centers their morality on the outcome of an action, rather than the action itself. Utilitarianism is a type of consequentialist theory, where an action that maximises utility, which is a metric that quantifies the outcome, is deemed morally praiseworthy. There are two types of utilitarians, act- and rule-utilitarians. While act-utilitarians scrutinise every act to ensure they maximise the overall well-being based on the utility principle, rule-utilitarians prioritise the rules which strictly maximise utility, irrespective of whether it is an optimal action. There are other variants of utilitarianism and consequentialist theories. However, for the purposes of this paper, which focuses on RL-based implementations, it would suffice to keep in mind that consequentialist ethics places a majority of the onus on the consequences.
Deontological ethics, as opposed to consequentialist theories, scrutinises an act based on whether it fulfilled a set of rules or duties. Deontology derives from Kantian ethics, who said, “So act that you use humanity, whether in your own person or in the person of any other, always at the same time as an end, never merely as a means.” [28]. There are several variants of deontology as well, and in 4.4 where we discuss our results, we delineate some of them based on the RL implementations. The third major theory is virtue ethics, which is a normative theory that centers on excellence of moral character. For example, temperance is regarded as a virtue where we seek a balance between overindulgence in pleasure and avoidance of pleasures. This balance is often referred to as the golden mean. Then there is prima-facie reasoning where duties such as fidelity, gratitude, etc., are central and decisions are made based on the context. This is different from deontology in the sense that duties in prima-facie reasoning are derived from human relations and experiences, rather than rational principles and categorical imperatives, as is the case in deontology. With this basic knowledge [^1] In the context of RL used in machine ethics, the moral theories are predominantly based on Western moral philosophy. However, we acknowledge that several there exist moral theories derived from Chinese, Indian, and African schools of thought, and it is plausible to implement them in reinforcement learning algorithms., we discuss next some of the systematic reviews in machine ethics.
2.3 Systematic reviews in machine ethics
@@
We briefly summarise the findings of the three reviews in machine ethics, their findings and methodology.
@@
Tolmeijer et al. [11] focused on the implementations of machine ethics specifically, and found 49 papers from Web of Science, Scopus, ACM Digital Library, Wiley Online Library, ScienceDirect, AAAI Publications, Springer Link, and IEEE Xplore. These papers were then analysed based on the types of ethical theory they implemented, the nontechnical aspects when implementing those theories, and technological details.
@@
Nallur [10] considers the landscape of machine ethics implementations, focusing on the different problems, techniques, methodologies and evaluation methodologies that are used. He did not do a systematic review, but rather a more qualitative analysis of selected prominent works.
@@
Yu et al. [12] survey advances in techniques for incorporating ethics into AI. They consider publications in “leading AI research conferences including AAAI, AAMAS, ECAI and IJCAI, as well as articles from well-known journals”. They organise the papers into four categories: Exploring Ethical Dilemmas, Individual Ethical Decision Frameworks, Collective Ethical Decision Frameworks; and 4. Ethics in Human-AI Interactions. There is, of course, a significant overlap among the papers included in all three surveys.
Cervantes et al. [29] performed a survey by first categorising artificial moral agents based on Moor’s definition [30]: implicit ethical agents, explicit ethical agents, and full ethical agents. Implicit agents are those whose actions are restricted to avoid unethical consequences, while explicit ethical agents follow ethical rules crafted explicitly for a purpose. Full ethical agents are those that are capable of reasoning and making ethical decisions like humans. Cervantes et al. analysed individual contributions in depth and highlighted strengths and weaknesses. Moreover, they categorised agents based on their approach to learning ethical principles: top-down (ethical principles pre-specified), bottom-up (infer ethical rules from their environment), and hybrid approaches (a combination of top-down and bottom-up). While not being a systematic literature review, Cervantes et al. highlighted features from examples based on a concise taxonomy.
@@
We here follow to a large degree the approach of Tolmeijer et al. [11] in terms of where we search for papers. This is because Tolmeijer et al. [11] is the only systematic review in machine ethics. Understanding the state of the art in machine ethics is not without challenges. The papers published in this area are sent to a wide spectrum of venues. Machine ethics is also not a very large area of research, and it is not very precisely positioned in the context of the AI alignment and AI Ethics initiatives. Therefore, it has to be accepted that any systematic review of the field is not likely to be exhaustive, even if we limit ourselves to a specific domain or methodology such as RL.
@@ 3 Survey methodology
In this section, we describe the research questions, the strategy used to find papers related to machine ethics and RL, following which we discuss the process used to review the found articles.
3.1 Research questions
Using the following formulated research questions (RQ), we plan the review procedure.
RQ1:
The first research question pertains to the RL algorithm and its corresponding components, such as the type of RL framework (MORL, DQN, etc.) adopted, the components of RL modified (state space, action space, policy function, value function, and reward function), and the RL environment (examples). Capturing these distinctions enables us to uncover trends in the field that researchers are inclined towards, for the development of artificial moral agents.
RQ2:
This question studies the type of ethical theories (deontology, consequentialism, etc.) realised by the RL agents and who decides the ethics. A strength of RL algorithms is to optimise an agent to achieve a certain goal. This research question offers insights into which ethical theories can be programmed and why researchers gravitate towards a certain theory. Moreover, understanding who decides what is ethical in a given context enables us to identify gaps in the field.
RQ3:
Which research contributions are prevalent in literature? Depending on how far a field has matured, it will be interesting to find out the proportion of purely theoretical, empirical and mathematical contributions.
We address RQ1, a portion of RQ2, and RQ3 in Section 4, and the remainder of RQ2 in 5.
3.2 Search strategy
Our review on implementations in machine ethics focuses on RL, hence we mandated this terminology while querying the databases. Since RL is to be used towards machine ethics implementations, we used alternative search terms based on seminal contributions in machine ethics, such as “artificial morality” [31, 29], “machine morality” and “robot ethics” [32], “computational ethics” [33], “roboethics” [34], “artificial moral agents” [29], “moral agent” [21], “value alignment” [25], “normative reasoning” [35], and “ethical dilemma” [36]. These terms have been used interchangeably to communicate the embedding of ethical theories into machines/computers/artificial intelligence agents. We show the query used below:
“reinforcement learning” AND ( “machine ethics” OR “artificial morality” OR “moral agent” OR “machine morality” OR “computational ethics” OR “roboethics” OR “robot ethics” OR “artificial moral agents” OR “value alignment” OR “normative reasoning” OR “ethical dilemma”)
The query was used to search the following databases (Table 1). In the right-hand side column we give the number of ‘hits’ from the query. Further filtering was performed on these ‘hits’.
@@ 3.2 Search strategy
ACM Digital Library 174
IEEE Xplore 54
ScienceDirect 78
Scopus 161
SpringerLink 251
Web of Science 28
Wiley Online Library 175
Total 921
@@
3.3 Selection process
A total of 921 papers were further filtered based on relevance to RL and machine ethics. We included only journal articles, conference papers, and book chapters which were published before 4th January 2024. The diagram in Figure 1 shows the elimination process we used to arrive at the final list of manuscripts.
Figure 1: The selection process used to eliminate irrelevant articles but keep the relevant ones. First is to check whether the article focuses on RL, followed by checking if the RL algorithm is used for moral decision-making, and finally to eliminate duplicates that appeared across multiple databases.
After applying these selection criteria, we ended up with 87 papers for review. Subsequently, we conducted a more focused review by filtering the papers based on their relevance to both machine ethics and RL, resulting in a total of 59 articles. Some examples of papers excluded from the 87 papers are [37, 26, 38]. For example, Hadfield-Menell et al. [26] specifically examined the issue of value alignment in conjunction with reinforcement learning methodologies, approaching the problem from a robotics-centric paradigm. The term “values”, as articulated by these researchers, pertains to the imperative for robotic systems to exhibit behaviours that are congruent with established human decision-making protocols, as opposed to adhering to a specific ethical framework or moral philosophy. On the other hand, Serramia et al. [38], while making “reinforcement learning” references, do not explicitly modify an existing RL algorithm to encode ethics. In the next section, we delve deeper into the taxonomy creation and review process.
4 The analysis
In this section, we discuss the taxonomy and categorisations used to organise our survey based on research questions, RQ1-RQ3 (except for a portion of RQ2 which resides in Discussion (Section 5).
4.1 Component of RL Modified
There are several aspects that affect RL algorithms and their performance on tasks. For instance, the challenge of designing a good reward function to correctly incentivise agents can make or break the agent’s performance. Other aspects such as policy functions, value functions, etc., also play a part. In terms of developing RL algorithms which make more ethical choices, several researchers modify either one or more of these components to achieve the desired result. We categorised the research articles based on which component of RL was modified, and we present our findings in Table 2.